OpenStudy (anonymous):

How many 3 digit positive even integers are there, such that all the digits are distinct from each other?

OpenStudy (raden):

case I : if the last digit be 0, then the position at the first number we can take 9 possible ways, and the 2nd position number there are 8 possible ways. it can represented be 9*8 = 72 ways case II : if the last digit is not 0, so there are 4 possible ways (2,4,6,8) for the last digit, then the position at the first we can take 8 possible ways,and the 2nd position there are 8 possible ways. it can represented be 8*8*4 = 256 ways. so, the total way = 72+256 = 328